A rectangle has a perimeter of 60 inches and length of 22 inches. What is the width of the rectangle?
Answer by rapaljer(4671)

Let x= width of the rectangle
2W + 2L = Perimeter
2x + 2(22) = 60
2x+44=60
2x=60-44
2x=16
x= 8 inches

Check: 2*8+2*22=60
16+44=60

For extra explanation on Perimeter problems, please see my Lesson Plans in algebra.com, or see also my own website by clicking on my tutor name "rapaljer" anywhere in algebra.com, look for the first link on my homepage which is "Basic and Intermediate Algebra: One Step at a Time", then look for Basic Algebra, Chapter 1, Section 1.10 Word Problems, Part 1.

R^2